OpenStudy (anonymous):

Solve the equation. –10 = 2k – 8

OpenStudy (texaschic101):

To solve this you have to get 2k on one side and everything else on the other side. So first, you add 8 to both sides.... -10 + 8 = 2k - 8 + 8 (combine like terms) - 2 = 2k (now you want to divide both sides by 2 to solve for k) -2/2 = 2k/2 -1 = k check... -10 = 2k - 8 -10 = 2(-1) - 8 -10 = - 2 - 8 - 10 = - 10 (correct) ANSWER : k = - 1